OpenStudy (anonymous):

Solve y = 2x - 4 by the substitution method. 7x – 5y = 14

OpenStudy (mathteacher1729):

Since you've already got " y equals ... something" in the first equation, substitute that into the second to solve for x. y = 2x - 4 7x – 5y = 14 Take the first equation sub it into the first to get... 7x – 5y = 14 7x – 5() = 14 7x – 5(2x - 4 ) = 14 can you solve that last one now?

OpenStudy (anonymous):

x=2 y=0

OpenStudy (mathteacher1729):

@Josh321 Usually instead of just giving the answer, we try to give students hints, and help them learn how to solve the problem themselves. :)

OpenStudy (anonymous):

thankss!

OpenStudy (anonymous):

fine 7x-5y=14 +5y +5y 7x=5y+14 -14 -14 7x-14=5y 7x-14=5(2x-4) 7x-14=10x-20 -7x -7x -14=3x-20 +20 +20 6=3x 6/3=3x/3 2=x y=2x-4 y=2(2)-4 y=4-4 y=0

OpenStudy (anonymous):

the type box messed up lol

OpenStudy (anonymous):

it's mushed

OpenStudy (mathteacher1729):

That is... quite a lengthy solution. y = 2x - 4 7x – 5y = 14 Take the first equation sub it into the first to get... 7x – 5y = 14 7x – 5() = 14 7x – 5(2x - 4 ) = 14 distribute the 5 on the left 7x -10x +20 = 14 subtract 20 from both sides, and combine 7x with -10x -3x = -6 divide by -3 on both sides x = 2 Now substitute x = 2 into the first equation y = 2x - 4 y = 2x - 4 y = 2(2) - 4 y = 4-4 y = 0 done x = 2 y = 4 and this agrees with the picture. @coolaidd not sure if this is helpful, but there it is anyway. :)